Consider first the goods market model with constant investment that we saw in Chapter 3. Consumption is given by:
C = Co + c1(Y-T)
And I, G and T are given.
a. Solve for equilibrium output. What is the value of the multiplier?
Now let investment depend on both sales and the interest rate:
I = b0 + b1Y -b2i
b. Solve for equilibrium output using the methods learned in chapter 3. At a given interest rate, why is the effect of a change in autonomous spending bigger than what it was in part (a)? Why? (Assume c1 + b1 = 1)
13
MAE206D /April 2020
c. Solve for equilibrium level of investment.
d. Let’s go behind the scene in the monetary market. Use the equilibrium in the money market M/P = d1Y – d2i to solve for the equilibrium level of the real money supply.
How does the real money supply vary with government spendin
1
Expert's answer
2020-04-22T11:41:14-0400

a.Y=[1/(1-c1)][c0-c1T+I+G], and the multiplier is equal to [1/(1-c1)]

b.Y=[1/(1-c1)][c0-c1T+b0 + b1Y -b2i +G]

the sensitivity of autonomous expenses to the interest rate (b) is high, which means that even a slight change in R leads to a significant change in autonomous expenses, as a result, of income

с.


"\frac{M}{P}=\frac{Y}{V}"

d1Y - d2i=(1/(1-c1)][c0-c1T+b0 + b1Y -b2i +G])/V


with increasing nominal government money supply increases, but this contributes to the acceleration of inflation, which leads to a decrease in real money supply